# Journey to Frontier-Level Performance of Foundation Models

### v1

@@ 1 Introduction
K-EXAONE builds on the hybrid architecture of EXAONE 4.0 [bae2026exaone40unifiedlarge], combining reasoning and non-reasoning capabilities to enhance both general-purpose and specialized use cases. It also uses a hybrid attention mechanism that integrates global and local attention modules, enabling efficient processing of long-context inputs—a critical feature for real-world applications.
@@ 2.1 Model Configurations
As illustrated in Figure 2, K-EXAONE employs a fine-grained sparse MoE design inspired by prior work [deepseekai2025deepseekv3technicalreport], consisting of 128 experts, where the top-8 experts are activated per token together with an additional shared expert, resulting in nine concurrently active experts per routing decision. Although the total number of parameters amounts to 236B, only approximately 23B parameters are activated, enabling high representational diversity and strong performance while maintaining resource-efficient training and inference. To improve routing stability and expert utilization efficiency, sequence-level load balancing is employed in the MoE routing mechanism, and a dropless routing policy [gale2023megablocks] is adopted to ensure that all tokens are dispatched to experts without capacity-based dropping, thereby stabilizing gradient flow and improving convergence behavior in large-scale MoE training.
@@
To enhance training stability and long-context extrapolation, K-EXAONE incorporates two architectural features—QK Norm and SWA (Sliding Window Attention) [beltagy2020longformer]-only RoPE (Rotary Positional Embeddings)—inherited from EXAONE 4.0. QK Norm applies layer normalization to the query and key vectors prior to attention computation, mitigating attention logit explosion in deep networks and stabilizing training dynamics, while RoPE are selectively applied only to SWA layers, preventing interference with global token interactions and improving robustness to long-sequence extrapolation.
@@ 2.2 Tokenizer
Compared to previous models in EXAONE series, we redesign the tokenizer and increase the vocabulary size from 100K to 150K to improve token efficiency, downstream task performance, and multilingual scalability. We retain the 70% high-frequency portion of the original vocabulary and reallocate capacity to expand coverage for additional languages, STEM (Science, Technology, Engineering, Mathematics), and code domains. To further improve efficiency, we adopt SuperBPE [liu2025superbpe] strategy that introduces superword tokens, allowing common word sequences to a single token and reducing overall sequence length. The superword tokens make up about 20% of the K-EXAONE’s vocabulary, allocated across English, Korean, and multilingual coverage in a 2:3:1 ratio.
@@
In addition, we update the pre-tokenization regex (regular expression) and normalization method to support the expanded vocabulary and superword units. We replace pre-tokenization regex for handling superword boundaries, line breaks, and multilingual Unicode characters. We also switch Unicode normalization [unicode-normalization-17] from NFKC to NFC to preserve semantic distinctions in superscripts, subscripts, and symbol-rich text commonly found in code and STEM corpora. Figure 3 shows tokenizer efficiency measured as bytes per token, where higher values indicate that each token represents a larger span of text in bytes. K-EXAONE consistently improves efficiency across English, Korean, multilingual, STEM, and code inputs, achieving an approximately 30% improvement on average over the previous EXAONE tokenizer.
@@ 3.1 Pre-training
We further enrich our data synthesis pipeline by extending the curation strategy of EXAONE 4.0 to incorporate explicit reasoning supervision. Motivated by [wang2025thinkingaugmentedpretraining], we generate document-grounded thinking trajectories and combine them with the source content into unified samples that encode step-by-step inference. These thinking-augmented corpora facilitate the transfer of reasoning behaviors and improve the effectiveness of subsequent post-training.
@@
K-EXAONE is natively trained with FP8 precision and achieves training loss curves comparable to those obtained under BF16 precision, demonstrating that FP8 training preserves optimization stability while enabling full quantization-aware convergence. We adopt the Muon optimizer [liu2025muon] for all training stages in conjunction with the Warmup–Stable–Decay (WSD) [dremov2025training] learning rate scheduler. The maximum learning rate is set to $3.0\times 10^{-4}$, with a linear warm-up phase followed by a stable plateau and a subsequent decay schedule. For MoE regularization, the sequence auxiliary loss coefficient is fixed to $1.0\times 10^{-4}$, and the expert bias update factor is also set to $1.0\times 10^{-4}$ throughout training. For the MTP objective, a loss weight of 0.05 is applied.
@@ 3.2 Context Length Extension
To ensure strong long-context performance, we place particular emphasis on a Long-Document Dataset during the extension phases, consisting of full-document sequences that can be consumed within a single training instance. We train on these samples in an end-to-end manner by feeding entire long-document sequences without truncation, encouraging the model to capture long-range dependencies. Stage 1 prioritizes stable performance up to 32K tokens, whereas Stage 2 increases the emphasis on long-document samples to better model dependencies up to 256K tokens. To systematically monitor potential performance degradation, we conduct (i) short-context evaluations following the same protocol used in pre-training, and (ii) Needle-In-A-Haystack (NIAH) [niah] tests to assess the model’s ability to retain and retrieve information from long contexts. Training is iteratively repeated until the model consistently achieves near-perfect NIAH performance across the target context ranges for each stage (“green light”), indicating that K-EXAONE successfully extends to 256K tokens without compromising overall performance.
@@ 3.3 Post-training
In K-EXAONE, the post-training process is primarily organized into three stages: (i) large-scale supervised fine-tuning (SFT) to learn to follow a variety of user instructions and produce corresponding responses, (ii) reinforcement learning (RL) on reasoning-intensive and verifiable tasks, and (iii) preference learning to align with human preferences. Most of the dataset generation pipelines follow those of EXAONE 4.0. To summarize, we first categorize instruction-following tasks into several domains and adopt distinct generation methods or experts. To enhance Korean-specific capabilities, we utilize public and institutional data provided by Korea Data Industry Promotion Agency (K-DATA) [^1] https://www.kdata.or.kr, which is a public institution under the Ministry of Science and ICT of South Korea. Raw data is filtered and converted into diverse datasets, such as DocQA and translation, and they are used for supervised fine-tuning.
@@ 3.4 Reinforcement Learning
To enhance the reasoning capability of the post-trained model, we perform reinforcement learning (RL) with verifiable rewards. We train in a multi-task setup covering math, code, STEM, and instruction-following. For verification, we use a combination of rule-based verifiers and an LLM-as-a-judge. For optimization, we use AGAPO [bae2026exaone40unifiedlarge] with an off-policy policy-gradient using truncated importance sampling.
For training efficiency at scale, we adopt an off-policy policy-gradient objective with truncated importance sampling, following prior work [minimax2025minimaxm1scalingtesttimecompute, khatri2025artscalingreinforcementlearning]. We also apply zero-variance filtering by dropping prompts whose sampled rollouts receive identical rewards, resulting in zero advantages. We employ both group-level advantage computation and global advantage normalization to capture both within-group relative reward signals and batch-level information. During training, we exclude the KL penalty to improve performance while avoiding unnecessary computation. Finally, we freeze the MoE router throughout RL training.
@@ 3.5 Preference Learning
After RL training, we perform a preference learning stage to better align the model with human preferences. In this stage, we aim to preserve reasoning performance while focusing training on general alignment domains such as chat, safety, instruction-following, agentic tool use, and creative writing. We propose GrouPER (Group-wise SimPER), an improved variant of SimPER [xiao2025simper], and show that it improves general-domain performance.
Inspired by the GRPO [shao2024deepseekmathpushinglimitsmathematical], GrouPER samples multiple responses for each query and trains the model using group-wise advantages. For each response, we compute a preference reward by combining rule-based rewards with rubric-based generative rewards that score responses along multiple dimensions. We then compute a group-level advantage from these scores and integrate it into the SimPER-style objective.
@@ 3.6 Data Compliance
Developing AI models requires a large amount of data, and the acquisition and utilization of this data can lead to various legal issues, such as copyright infringement, intellectual property infringement, and personal information protection violations. To minimize these risks, LG AI Research conducts AI Compliance reviews throughout the entire process of data collection, AI model training, and information provision. For more detailed information, please refer to the EXAONE 3.0 Technical Report [an2026exaone3078binstruction] and the LG AI Ethics Principles [lgethics].
@@ 4.1 Benchmarks and Setup
  • World Knowledge: MMLU-Pro [wang2024mmlupro], GPQA-Diamond [rein2024gpqa], and Humanity’s Last Exam [^2] We use the text-only subset. [phan2025humanitysexam]
  • Math: IMO-AnswerBench [luong-etal-2025-towards], AIME 2025 [maa2025aime], and HMMT Nov 2025 [balunovic_srimatharena_2025]
  • Coding / Agentic Coding: LiveCodeBench Pro (25Q2 Medium) [zheng2025livecodebench], LiveCodeBench v6 [jain2025livecodebench],
    Terminal-Bench 2.0 [tbench_2025], and SWE-bench Verified [swebenchverified]
  • Agentic Tool Use: $\tau^{2}$-Bench [barres2025tau2benchevaluatingconversationalagents] and BrowseComp [wei2025browsecompsimplechallengingbenchmark]
  • Instruction Following: IFBench [pyatkin2025generalizing] and IFEval [zhou2023instructionfollowingevaluationlargelanguage]
  • Long Context Understanding: AA-LCR [artificialanalysis2025lcr] and OpenAI-MRCR [openai2025mrcr_snapshot]
  • Korean: KMMLU-Pro [hong-etal-2025-kmmlu], KoBALT [shin2025kobaltkoreanbenchmarkadvanced], CLIcK [kim-etal-2024-click], HRM8K [ko-etal-2025-understand], and Ko-LongBench (in-house)
  • Multilinguality [^3] We only evaluate five non-English supported languages on multilingual benchmarks: Korean (ko), German (de), Spanish (es), Japanese (ja), and Vietnamese (vi).: MMMLU [hendrycks2021measuring] and WMT24++ [deutsch-etal-2025-wmt24]
  • Safety: WildJailbreak [jiang2024wildteaming] and KGC-Safety [^4] Korean Global Civic Safety Benchmark. See Appendix F for details. (in-house)
To evaluate our model, we set the temperature to 1.0 and top-p [Holtzman2020The] to 0.95. We set the context length to 160K for the Long Context Understanding benchmarks, while 128K for others. We disable the MTP at inference time. For baseline models, when official scores are unavailable, we evaluate them in our internal environment with inference parameters set to the recommended configuration for each model. Please refer to Appendix C for the detailed evaluation setup of each benchmark and Appendix D for the in-house benchmarks.
@@ C.5 SWE-Bench Verified
We evaluate our model using mini-SWE-agent [yang2024sweagent] as our default agent, and use the same setup to obtain scores when official results are unavailable.
@@ C.9 WMT24++
Figure 8 presents the judging prompt for WMT24++. We adopt the judge prompt from official implementation [deutsch-etal-2025-wmt24], but we use gpt-5-mini-2025-08-07 as a judge model. The final scores are average translation scores from LLM judge between en $\leftrightarrow$ five non-English supported languages.
@@ C.10 WildJailbreak
WildJailbreak results by jointly analyzing ⟨input–model output⟩ pairs using the Qwen3Guard-Gen-8B model [zhao2025qwen3guard] to determine whether responses are safe. Performance is reported using the Safe Rate, defined as the proportion of test cases classified as safe across the full test set, where a higher Safe Rate indicates a safer model.
@@ D.1 Code Utility Benchmark (CodeUtilityBench)
CodeUtilityBench consists of 300 test instances, with 20 instances allocated to each of the 15 tasks (excluding the visualize task). Each test instance includes five evaluation rubrics, which are generated following task-specific protocols. To ensure dataset quality, human experts reviewed all queries and rubrics and replaced unsuitable items. For the visualize task, we use AutoCodeArena [zhuo2025bigcodearena], excluding instances from the Problem Solving category as they are not applicable to our task definition.
@@ D.2 Ko-LongBench
Ko-LongBench is an in-house benchmark designed to evaluate long-context understanding in Korean. It comprises a diverse set of tasks including Document QA, Story Understanding, Dialogue History Understanding, In-Context Learning, Structured QA, and RAG to assess LLMs’ long-context capabilities in practical settings. Detailed benchmark descriptions, dataset statistics, and representative prompt examples are available in the EXAONE 4.0 Technical Report [bae2026exaone40unifiedlarge].

### v2

@@ 1 Introduction
K-EXAONE builds on the hybrid architecture of EXAONE 4.0 [2], combining reasoning and non-reasoning capabilities to enhance both general-purpose and specialized use cases. It also uses a hybrid attention mechanism that integrates global and local attention modules, enabling efficient processing of long-context inputs—a critical feature for real-world applications.
@@ 2.1 Model Configurations
As illustrated in Figure 2, K-EXAONE employs a fine-grained sparse MoE design inspired by prior work [7], consisting of 128 experts, where the top-8 experts are activated per token together with an additional shared expert, resulting in nine concurrently active experts per routing decision. Although the total number of parameters amounts to 236B, only approximately 23B parameters are activated, enabling high representational diversity and strong performance while maintaining resource-efficient training and inference. To improve routing stability and expert utilization efficiency, sequence-level load balancing is employed in the MoE routing mechanism, and a dropless routing policy [10] is adopted to ensure that all tokens are dispatched to experts without capacity-based dropping, thereby stabilizing gradient flow and improving convergence behavior in large-scale MoE training.
@@
To enhance training stability and long-context extrapolation, K-EXAONE incorporates two architectural features—QK Norm and SWA (Sliding Window Attention) [5]-only RoPE (Rotary Positional Embeddings)—inherited from EXAONE 4.0. QK Norm applies layer normalization to the query and key vectors prior to attention computation, mitigating attention logit explosion in deep networks and stabilizing training dynamics, while RoPE are selectively applied only to SWA layers, preventing interference with global token interactions and improving robustness to long-sequence extrapolation.
@@ 2.2 Tokenizer
Compared to previous models in EXAONE series, we redesign the tokenizer and increase the vocabulary size from 100K to 150K to improve token efficiency, downstream task performance, and multilingual scalability. We retain the 70% high-frequency portion of the original vocabulary and reallocate capacity to expand coverage for additional languages, STEM (Science, Technology, Engineering, Mathematics), and code domains. To further improve efficiency, we adopt SuperBPE [22] strategy that introduces superword tokens, allowing common word sequences to a single token and reducing overall sequence length. The superword tokens make up about 20% of the K-EXAONE’s vocabulary, allocated across English, Korean, and multilingual coverage in a 2:3:1 ratio.
@@
In addition, we update the pre-tokenization regex (regular expression) and normalization method to support the expanded vocabulary and superword units. We replace pre-tokenization regex for handling superword boundaries, line breaks, and multilingual Unicode characters. We also switch Unicode normalization [35] from NFKC to NFC to preserve semantic distinctions in superscripts, subscripts, and symbol-rich text commonly found in code and STEM corpora. Figure 3 shows tokenizer efficiency measured as bytes per token, where higher values indicate that each token represents a larger span of text in bytes. K-EXAONE consistently improves efficiency across English, Korean, multilingual, STEM, and code inputs, achieving an approximately 30% improvement on average over the previous EXAONE tokenizer.
@@ 3.1 Pre-training
We further enrich our data synthesis pipeline by extending the curation strategy of EXAONE 4.0 to incorporate explicit reasoning supervision. Motivated by [36], we generate document-grounded thinking trajectories and combine them with the source content into unified samples that encode step-by-step inference. These thinking-augmented corpora facilitate the transfer of reasoning behaviors and improve the effectiveness of subsequent post-training.
@@
K-EXAONE is natively trained with FP8 precision and achieves training loss curves comparable to those obtained under BF16 precision, demonstrating that FP8 training preserves optimization stability while enabling full quantization-aware convergence. We adopt the Muon optimizer [23] for all training stages in conjunction with the Warmup–Stable–Decay (WSD) [9] learning rate scheduler. The maximum learning rate is set to $3.0\times 10^{-4}$, with a linear warm-up phase followed by a stable plateau and a subsequent decay schedule. For MoE regularization, the sequence auxiliary loss coefficient is fixed to $1.0\times 10^{-4}$, and the expert bias update factor is also set to $1.0\times 10^{-4}$ throughout training. For the MTP objective, a loss weight of 0.05 is applied.
@@ 3.2 Context Length Extension
To ensure strong long-context performance, we place particular emphasis on a Long-Document Dataset during the extension phases, consisting of full-document sequences that can be consumed within a single training instance. We train on these samples in an end-to-end manner by feeding entire long-document sequences without truncation, encouraging the model to capture long-range dependencies. Stage 1 prioritizes stable performance up to 32K tokens, whereas Stage 2 increases the emphasis on long-document samples to better model dependencies up to 256K tokens. To systematically monitor potential performance degradation, we conduct (i) short-context evaluations following the same protocol used in pre-training, and (ii) Needle-In-A-Haystack (NIAH) [12] tests to assess the model’s ability to retain and retrieve information from long contexts. Training is iteratively repeated until the model consistently achieves near-perfect NIAH performance across the target context ranges for each stage (“green light”), indicating that K-EXAONE successfully extends to 256K tokens without compromising overall performance.
@@ 3.3 Post-training
In K-EXAONE, the post-training process is primarily organized into three stages: (i) large-scale supervised fine-tuning (SFT) to learn to follow a variety of user instructions and produce corresponding responses, (ii) reinforcement learning (RL) on reasoning-intensive and verifiable tasks, and (iii) preference learning to align with human preferences. Most of the dataset generation pipelines follow those of EXAONE 4.0. To summarize, we first categorize instruction-following tasks into several domains and adopt distinct generation methods or experts. To enhance Korean-specific capabilities, we utilize public and institutional data provided by Ministry of Science and ICT (MSIT) of South Korea and its affiliated agencies, such as National Information Society Agency (NIA) [^1] https://www.nia.or.kr and Korea Data Industry Promotion Agency (K-DATA) [^2] https://www.kdata.or.kr.
@@ 3.4 Reinforcement Learning
To enhance the reasoning capability of the post-trained model, we perform reinforcement learning (RL) with verifiable rewards. We train in a multi-task setup covering math, code, STEM, and instruction-following. For verification, we use a combination of rule-based verifiers and an LLM-as-a-judge. For optimization, we use AGAPO [2] with an off-policy policy-gradient using truncated importance sampling.
For training efficiency at scale, we adopt an off-policy policy-gradient objective with truncated importance sampling, following prior work [26, 18]. We also apply zero-variance filtering by dropping prompts whose sampled rollouts receive identical rewards, resulting in zero advantages. We employ both group-level advantage computation and global advantage normalization to capture both within-group relative reward signals and batch-level information. During training, we exclude the KL penalty to improve performance while avoiding unnecessary computation. Finally, we freeze the MoE router throughout RL training.
@@ 3.5 Preference Learning
After RL training, we perform a preference learning stage to better align the model with human preferences. In this stage, we aim to preserve reasoning performance while focusing training on general alignment domains such as chat, safety, instruction-following, agentic tool use, and creative writing. We propose GrouPER (Group-wise SimPER), an improved variant of SimPER [39], and show that it improves general-domain performance.
Inspired by the GRPO [31], GrouPER samples multiple responses for each query and trains the model using group-wise advantages. For each response, we compute a preference reward by combining rule-based rewards with rubric-based generative rewards that score responses along multiple dimensions. We then compute a group-level advantage from these scores and integrate it into the SimPER-style objective.
@@ 3.6 Data Compliance
Developing AI models requires a large amount of data, and the acquisition and utilization of this data can lead to various legal issues, such as copyright infringement, intellectual property infringement, and personal information protection violations. To minimize these risks, LG AI Research conducts AI Compliance reviews throughout the entire process of data collection, AI model training, and information provision. For more detailed information, please refer to the EXAONE 3.0 Technical Report [1] and the LG AI Ethics Principles [21].
@@ 4.1 Benchmarks and Setup
  • World Knowledge: MMLU-Pro [37], GPQA-Diamond [30], and Humanity’s Last Exam [^3] We use the text-only subset. [28]
  • Math: IMO-AnswerBench [24], AIME 2025 [25], and HMMT Nov 2025 [3]
  • Coding / Agentic Coding: LiveCodeBench Pro (25Q2 Medium) [42], LiveCodeBench v6 [16],
    Terminal-Bench 2.0 [34], and SWE-bench Verified [6]
  • Agentic Tool Use: $\tau^{2}$-Bench [4] and BrowseComp [38]
  • Instruction Following: IFBench [29] and IFEval [43]
  • Long Context Understanding: AA-LCR [33] and OpenAI-MRCR [27]
  • Korean: KMMLU-Pro [15], KoBALT [32], CLIcK [19], HRM8K [20], and Ko-LongBench (in-house)
  • Multilinguality [^4] We only evaluate five non-English supported languages on multilingual benchmarks: Korean (ko), German (de), Spanish (es), Japanese (ja), and Vietnamese (vi).: MMMLU [13] and WMT24++ [8]
  • Safety: WildJailbreak [17] and KGC-Safety [^5] Korean Global Civic Safety Benchmark. See Appendix F for details. (in-house)
To evaluate our model, we set the temperature to 1.0 and top-p [14] to 0.95. We set the context length to 160K for the Long Context Understanding benchmarks, while 128K for others. We disable the MTP at inference time. For baseline models, when official scores are unavailable, we evaluate them in our internal environment with inference parameters set to the recommended configuration for each model. Please refer to Appendix C for the detailed evaluation setup of each benchmark and Appendix D for the in-house benchmarks.
@@ C.5 SWE-Bench Verified
We evaluate our model using mini-SWE-agent [40] as our default agent, and use the same setup to obtain scores when official results are unavailable.
@@ C.9 WMT24++
Figure 8 presents the judging prompt for WMT24++. We adopt the judge prompt from official implementation [8], but we use gpt-5-mini-2025-08-07 as a judge model. The final scores are average translation scores from LLM judge between en $\leftrightarrow$ five non-English supported languages.
@@ C.10 WildJailbreak
WildJailbreak results by jointly analyzing ⟨input–model output⟩ pairs using the Qwen3Guard-Gen-8B model [41] to determine whether responses are safe. Performance is reported using the Safe Rate, defined as the proportion of test cases classified as safe across the full test set, where a higher Safe Rate indicates a safer model.
@@ D.1 Code Utility Benchmark (CodeUtilityBench)
CodeUtilityBench consists of 300 test instances, with 20 instances allocated to each of the 15 tasks (excluding the visualize task). Each test instance includes five evaluation rubrics, which are generated following task-specific protocols. To ensure dataset quality, human experts reviewed all queries and rubrics and replaced unsuitable items. For the visualize task, we use AutoCodeArena [44], excluding instances from the Problem Solving category as they are not applicable to our task definition.
@@ D.2 Ko-LongBench
Ko-LongBench is an in-house benchmark designed to evaluate long-context understanding in Korean. It comprises a diverse set of tasks including Document QA, Story Understanding, Dialogue History Understanding, In-Context Learning, Structured QA, and RAG to assess LLMs’ long-context capabilities in practical settings. Detailed benchmark descriptions, dataset statistics, and representative prompt examples are available in the EXAONE 4.0 Technical Report [2].